# The impact of local noise recorded at the ET candidate sites on the signal to noise ratio of CBC gravitational wave signals for the ET triangle configuration

### v5

@@ Abstract
I Introduction
@@
For what concerns the configuration of ET, there are two proposals under consideration. The most recent one is that of a detector network composed by two widely separated L-shaped detectors with 15 km long arms[8, 16]. On the other hand, the original project foresees three pairs of nested interferometers arranged in an equilateral triangle (also called xylophone configuration, Figure 1) with the sides 10 km long [12, 13, 14, 15]. For each interferometer pair, one detector is optimized for low frequencies (2 Hz $<f<$ 40 Hz) and the other for high frequencies ($f>$ 40 Hz). Since recently, moving the lower limit to 3 Hz is being considered as well [15], but, to be on the conservative side, we keep using the 2 Hz lower limit for this work. In both the 2L and the triangle configurations, ET will be hosted underground, at a currently planned depth between 200 m and 300 m to reduce seismic motion at the input of the suspension system of the mirrors and to reduce the impact of atmospheric disturbances [17] and Newtonian noise (NN). In this work, since it poses the most challenges from the point of view of environmental noise mitigation, we focus on the ET triangular configuration.
@@
Generally speaking, the extension of the bandwidth to 2 Hz and the sharp increase in sensitivity will significantly improve the rate of detected events giving the possibility to issue early warnings for the coalescence of compact objects (CBC) several minutes, if not hours depending on the source, before the merger [18, 7, 19, 8, 20]. In fact, with respect to current detectors, CBC signals will spend more time in the ET accessible bandwidth, therefore enabling early detection and sky localization. As it was demonstrated by GW170817 [21, 22, 23, 24], this is extremely important in the case of binary neutron star (BNS) mergers. As a matter of fact, joint multimessenger observations with electromagnetic observatories are paramount to exploit the scientific potential of BNS events and shed new light on the internal structure of neutron stars [7]. In particular, depending on the distance of the source and requiring a signal-to-noise ratio (SNR) $\geq 12$ and a sky localization smaller than $100\rm\,deg^{2}$, ET is expected to be able to send early warnings between 1 and 20 hours before the BNS merger [7]. The release of an early warning increases the chances of detection for the electromagnetic counterparts. Astronomers would be able to point the telescope in the region of the signal to start the monitoring to obtain pre-merger images as well as merger. This enables to detect the early electromagnetic emission, which is fundamental to understand the physics of the emission mechanism and the merger remnant [7, 25].
@@
The significant increase in sensitivity below $100\rm\,Hz$ will also allow for more detailed observations about the merger of intermediate mass black holes (IMBH) [26, 27, 7, 28], i.e. black holes (BH) with masses in the $[10^{2},10^{5}]\rm\,M_{\bigodot}$ range. Their typical merger frequency is $<100\rm\,Hz$ and the signals are expected to spend a maximum of $\mathcal{O}(100\rm\,s)$ in the ET bandwidth. Currently, GW detectors are limited to BH masses $\mathcal{O}(100\rm\,M_{\bigodot})$ and hypothetical IMBH merger signals are typically only detectable in very short time intervals of a few milliseconds. On the other hand, ET will open the possibility of detecting these IMBH, studying the possibility that they are the seeds of the supermassive BHs in the center of galaxies [7, 28]. Furthermore, a precise reconstruction of waveforms for compact objects during inspiral and merger will allow accurate tests of General Relativity.
@@
Figure 1: Scheme of the ET xylophone for the triangular configuration. Each detector ET1,2,3 is composed of an interferometer optimized for low-frequency detection and another optimized for high-frequency detection for a total of six.
Since 2014, the Sos Enattos [56, 57, 58, 60, 59, 61] and EMR [62, 63, 64] sites have been the subject of characterization studies aimed at assessing their suitability to host ET. However, except [60] and the recent work published by [65], no works are assessing the impact of site dependent noise on the detectability of GW sources. Therefore, the scope of this paper is to provide a first estimation of how site dependent low-frequency noise may affect the SNR of CBC signals after modifying the ET sensitivity curve according to the actual seismic noise measured at each of the aforementioned ET candidate sites. We think that this approach may provide some relevant information for both the site selection process and the design of noise suppression systems for ET. In particular, extending the methodology already shown in [60] and [66], we investigate the case of IMBH and BNS focusing on the $[2,10]\rm\,Hz$ range, where ambient noise is more prominent [58, 59, 60] and may affect early warnings. Beyond 10 Hz, noise contributions from the detector hardware dominate the noise budget and, therefore, are out of the scope of our work. As a consequence of the focus in the low frequency range, our goal is not to claim whether a merger will be detected or not (merger frequencies are way beyond the considered frequency range), but only if it will be possible to observe the inspiral preceding the merger for multimessenger purposes. On the other hand, the details of how parameter estimation and other cosmological implications would be affected are beyond the goal of this work and will not be discussed here, but may represent one of the future developments of this study.
The paper is organized in the following way: in Section II we discuss the astrophysical background of this study; the method is presented in Section III; Section IV talks through the data used for this study; the results are discussed in Section V and VI, whereas Section VII gives a qualitative contribution to the discussion regarding the triangle and 2L configurations. Finally, the conclusions are debated in Section VIII.
@@
II Astrophysical background
For a long time, scientists questioned about the existence of a class of black holes, identified as IMBH [27, 26, 28], between those formed by stellar collapse (i.e. $M<10^{2}\rm M_{\bigodot}$) and the supermassive BH at the centers of galaxies (i.e. $M>10^{5}\rm M_{\bigodot}$). Before the third observing run (O3) [70] of aLIGO and AdVirgo, there was no clear evidence about their existence and the gap between $[10^{2},10^{5}]\rm M_{\bigodot}$ remained unexplained. During the third observing run (O3) of the LIGO-Virgo detectors, GW190521 [71], consistent with two BH merging to form a $\simeq 140\rm M_{\bigodot}$ remnant, provided the first direct evidence of IMBH formation. During the same observing run, a few more marginal candidates for IMBH formation were found, but none sufficiently significant to indicate detection of further IMBH mergers [70, 72].
The formation channel of IMBH has not been completely established. A possible explanation is that they may have formed via the subsequent, usually referred to as hierarchical, mergers of lighter BH [73]. Other IMBH formation channels also include the direct collapse of massive first-generation, low-metallicity Population III stars [28]. The existence of IMBH and their accretion through hierarchical mergers may also provide relevant information about the formation of supermassive BH [70].
Since the merger frequency of two compact objects orbiting around each other is inversely proportional to their chirp mass $M_{c}\equiv\frac{(m_{1}m_{2})^{3/5}}{(m_{1}+m_{2})^{1/5}}$ as $f_{GW}\propto M_{c}^{-5/8}$, a BH binary with component masses in the IMBH range merge at relatively low frequencies, i.e. $f<100$ Hz. For example, using the Keplerian approximation for circular orbits and considering the two BH as merged when their orbital separation is less than the Schwarzschild radius
$R_{s}=\frac{2Gm}{c^{3}},$ (1)
an equal mass IMBH with $m_{1}=m_{2}=100\rm M_{\bigodot}$, each with $R_{s}\simeq 533\rm km$, has an orbital frequency at merger [74]
$\omega_{k}(R_{s})=\sqrt{\frac{Gm_{\rm{tot}}}{R_{s}^{3}}}\simeq 420\rm Hz,$ (2)
which gives an approximate merger frequency of
$f_{GW}=\frac{\omega_{GW}}{2\pi}=\frac{\omega_{k}}{\pi}\simeq 133\rm\,Hz.$ (3)
Furthermore, if we calculate the number of cycles spent in a given frequency band $[f_{\rm{min}},f_{\rm{max}}]$ [74]
$\mathcal{N}_{\rm{cyc}}=\frac{1}{32\pi^{8/3}}\left(\frac{GM_{c}}{c^{3}}\right)^{-5/3}(f_{\rm{min}}^{-5/3}-f_{\rm{max}}^{-5/3}),$ (4)
we find that most of the cycles of the observable waveform are contained in $[2,10]\rm\,Hz$. At cosmological distances within the reach of ET [7], redshift $z$ also contributes to lowering the observed merger frequency as masses observed on Earth in the detector frame have to be corrected as $m_{\rm{det}}=m_{\rm{sour}}(1+z)$.
((a)) Time-frequency evolution of an equal mass IMBH with $m_{\rm{tot}}=200\rm\,M{\bigodot}$ in the circular orbit, non spinning case simulated using the IMRPhenomD waveform approximant [75, 76].
((b)) Time-frequency evolution of an equal mass BNS with $m_{\rm{tot}}=2.8\rm\,M{\bigodot}$ in the circular orbit, non spinning case simulated using the IMRPhenomPv2_NRTidalv2 waveform approximant [77].
Figure 3: Examples of time-frequency evolution of CBC signals.
Figure 3(a) shows the time-frequency evolution of an equal mass $m_{tot}=200\rm\,M_{\bigodot}$ IMBH in the circular orbit, non spinning case computed using the IMRPhenomD waveform approximant [75, 76]. In this case, the merger frequency appears to be $f_{merge}\simeq 82\rm\,Hz$, lower than what is shown in Equation 3. This is due to the fact that GW waveform approximants take into account relativistic effects close to the merger. In fact, we refer to the innermost stable circular orbit (ISCO) as the last orbit beyond which stable circular orbits are no longer allowed. For the case mentioned above, $f_{ISCO}\simeq 11\rm Hz$. In particular, IMRPhenomD is based on a combination of analytic post-Newtonian (PN) and effective-one-body (EOB) methods to describe the inspiral, whereas the merger phase is calibrated using numerical relativity models [75, 76]. Figure 3(a) also highlights the importance of the $[2,10]\rm\,Hz$ band, in which an IMBH spends most of the time before the merger.
BNS, on the other hand, are a completely different source. They are composed of two NS, have an expected total mass between $2.2\rm M_{\bigodot}$ and $4.4\rm M_{\bigodot}$ and merge at frequency $\mathcal{O}(\rm kHz)$. All the relevant information carried by GW about the physics of neutron stars can be found in the high-frequency part of the signal as well. However, low-frequency detection plays a crucial role in the observation of BNS mergers. In fact, as mentioned in Section I, detecting a BNS as soon as possible may not only significantly increase the observation rate [19, 8, 78], but also allow for precise sky localization [20] and sending prompt early warnings to observatories for conducting multimessengers observing campaigns [18, 7]. Moreover, BNS typically spend several hours at low frequency accumulating SNR before the chirp. Figure 3(b) shows the time-frequency evolution of an equal mass $m_{\rm{tot}}=2.8\rm\,M_{\bigodot}$ BNS underlining the fact that it spends $\simeq 20\rm\,hours$ in the $[2,10]\rm\,Hz$ band. Furthremore, Equation 4 shows that $\mathcal{N}_{\rm{cyc}}([2,10]\rm\,Hz])\simeq 7\times 10^{5}$ and $\mathcal{N}_{\rm{cyc}}([2,4096]\rm\,Hz])\simeq 7.6\times 10^{5}$. This means that only fully accomplishing the low-frequency design sensitivity, especially below 10 Hz, may allow to exploit the potential of ET for what concerns multimessenger astronomy [8, 19].
III Methodology
@@
To estimate the impact of site-dependent noise over the SNR of CBC events, we first calculate how site noise changes the ET design sensitivity curve. For this study we use the design curve already applied for [8]. Depending on the frequency band, different contributions affect the ET noise budget curve like seismic, Newtonian, thermal, and quantum noise. Among these noise sources, the dominant contribution in the frequency region between 2 and 10 Hz is the NN. Beyond 10 Hz, other contributions, like quantum and thermal noise from the detector hardware, dominate the noise budget [79, 80, 54, 81]. NN is caused by gravity fluctuations from mass distributions that change around the detector with time. These fluctuations are generated by variations of atmospheric and ground densities in the vicinity of the test masses of the detectors. Ground density variations are caused by ground motion, i.e., seismic waves, which are the only contribution to NN considered in this work. Atmospheric NN is not taken into account due to the current lack of dedicated environmental sensors installed at the ET candidate sites.
@@
which is Equation 5 of [60] and Equation 6 of [81], where $\vec{\xi}_{P,S}$ is the seismic displacement vector for the compressional (P) and shear (S) waves, respectively, and $\omega$ is the angular frequency. Equation 5 holds for a spherical cavern in an infinite homogeneous medium as it is clearly stated after Equation 6 of [81]. An even more detailed derivation can be found in Section 3.3 of [54]. In general, equation 5 is evaluated assuming:
@@
Referring to equation 5, the used ASD $\tilde{x}$ are the percentiles (10th, 50th, and 90th) of the spectral distributions calculated on $120\rm\,s$ windows. The length of the time window is chosen according to the typical maximum duration of an IMBH signal (Figure 5(b)) and on the time segment duration proposed for CBC multiband analysis for BNS [20]. Under these assumptions, the ET noise budget that we use takes into account the NN evaluated from the seismic noise of each candidate site for the 10th, 50th, and 90th percentile of seismic spectra. Since a simple projection alone does not show the effect of non-stationarity of NN (from now on defined glitchness) we also use a more effective indicator, called the Noise to Target Ratio (NTR) and derived from [60] as:
[…]
@@
After inferring the modified ET sensitivity curves, we gather the events of interest for our study from astrophysical catalogs and we use PyCBC [82] to generate and inject signals in simulated ET noise. Waveforms are generated using the IMRPhenomD [75, 76] GW approximant for IMBH and IMRPhenomPv2_NRTidalv2 [77] for BNS, both available in the LIGO-Virgo-KAGRA Algorithm Library (LAL) [83]. To take into account the antenna pattern of the detector, the ET triangle is simulated considering three properly oriented L-shaped co-located detectors, with an arm length of $10\rm\,km$ and an arm angle of $60^{\circ}$. The variation of the antenna pattern with time is also taken into account. Since the actual orientation of the ET triangle has not been decided yet, we choose an arbitrary positioning of the detector. For the Sardinia candidate site, ET is assumed to be located at the Sos Enattos mine at $[40.44^{\circ}\rm\,N,9.44^{\circ}\rm\,E]$, whereas for the EMR site, we have chosen the village of Terziet at $[50.73^{\circ}\rm\,N,5.91^{\circ}\rm\,E]$ as a reference. For each detector, noise is obtained by generating white noise colored according to the given sensitivity curve and the same sources are taken into account. Waveforms are injected into the noise of each detector composing ET (labeled ET1, ET2 and ET3 respectively).
@@
where $\tilde{s}(f)$ is the Fourier Transform (FT) of the data, $\tilde{h}^{*}(f)$ is the FT of the template and $h_{cos}$/$h_{sin}$ are the two orthogonal phase components of the template shown in Equations 9 and 10 [84, 85]. In our case, $f_{\rm{min}}=2\rm\,Hz$ and $f_{\rm{max}}=10\rm\,Hz$ because, as mentioned above, we consider ambient noise only which modifies the ET sensitivity curve not beyond the $[2,10]\rm\,Hz$ range. Moreover, taking into account Equations 11 and 12 we note that MSNR out of this bandwidth gives the same contribution to the total MSNR in all cases and would not bring relevant information. Therefore we restrict the calculation between 2 Hz and 10 Hz only. We also point out that we assume perfect knowledge of the source, i.e., matched filter is calculated using the same waveform template injected in the data.
IV Data
IV.1 Seismic

### v6

@@ Abstract
Keywords:
Einstein Telescope, Gravitational waves, Newtonian noise, compact binary coalescence
1 Introduction
@@
For what concerns the configuration of ET, there are two proposals under consideration. The most recent one is that of a detector network composed by two widely separated L-shaped detectors with 15 km long arms[8, 16]. On the other hand, the original project foresees three pairs of nested interferometers arranged in an equilateral triangle (Figure 1) with the sides 10 km long [12, 13, 14, 15]. For each interferometer pair, one detector is optimized for low frequencies (2 Hz $<f<$ 40 Hz) and the other for high frequencies ($f>$ 40 Hz). Since recently, moving the lower limit to 3 Hz is being considered as well [15], but, to be on the conservative side, we keep using the 2 Hz lower limit for this work. In both the 2L and the triangle configurations, ET will be hosted underground, at a currently planned depth between 200 m and 300 m to reduce seismic motion at the input of the suspension system of the mirrors and to reduce the impact of atmospheric disturbances [17] and Newtonian noise (NN). In this work, since it poses the most challenges from the point of view of environmental noise mitigation, we focus on the ET triangular configuration.
@@
Generally speaking, the extension of the bandwidth to 2 Hz and the sharp increase in sensitivity will significantly improve the rate of detected events giving the possibility to issue early warnings for the coalescence of compact objects (CBC) several minutes, if not hours depending on the source, before the merger [18, 7, 19, 8, 20, 21]. In fact, with respect to current detectors, CBC signals will spend more time in the ET accessible bandwidth, therefore enabling early detection and sky localization. As it was demonstrated by GW170817 [22, 23, 24, 25, 21], this is extremely important in the case of binary neutron star (BNS) mergers. As a matter of fact, joint multimessenger observations with electromagnetic observatories are paramount to exploit the scientific potential of BNS events and shed new light on the internal structure of neutron stars [7]. In particular, depending on the distance of the source and requiring a signal-to-noise ratio (SNR) $\geq 12$ and a sky localization smaller than $100\rm\,deg^{2}$, ET is expected to be able to send early warnings between 1 and 20 hours before the BNS merger [7, 21]. The release of an early warning increases the chances of detection for the electromagnetic counterparts. Astronomers would be able to point the telescope in the region of the signal to start the monitoring to obtain pre-merger images, ensuring the coverage of the merger and post merger phases as well. This enables to detect the early electromagnetic emission, which is fundamental to understand the physics of the emission mechanism and the merger remnant [7, 21]. Moreover, this class of signals will spend most of their time in the ET bandwidth in the $[2,10]\rm\,Hz$ range, thus making the low frequency sensitivity of ET extremely important for early detection.
@@
The significant increase in sensitivity below $100\rm\,Hz$ will also allow for more detailed observations about the merger of intermediate mass black holes (IMBH) [26, 27, 7, 28], i.e. black holes (BH) with masses in the $[10^{2},10^{5}]\rm\,M_{\bigodot}$ range, of which there was no clear evidence about their existence. This until the third observing run (O3) of the LIGO-Virgo detectors, which observed an event consistent with two BH merging to form a $\simeq 140\rm M_{\bigodot}$ remnant, providing the first direct evidence of IMBH formation [29]. During the same observing run, a few more marginal candidates for IMBH formation were found, but none sufficiently significant to indicate detection of further IMBH mergers [30, 31]. The typical merger frequency of this class of astrophysical objects is $<100\rm\,Hz$ and the signals are expected to spend a maximum of $\mathcal{O}(100\rm\,s)$ in the ET bandwidth with most cycles of their waveform contained in the $[2,10]\rm\,Hz$ band. Currently, GW detectors are limited to BH masses $\mathcal{O}(100\rm\,M_{\bigodot})$ and IMBH merger signals are typically only detectable in very short time intervals of a few milliseconds. On the other hand, ET will open the possibility of detecting these IMBH, studying their formation channels and the possibility that they are the seeds of the supermassive BHs in the center of galaxies [7, 28]. Furthermore, a precise reconstruction of waveforms for compact objects during inspiral and merger will allow accurate tests of General Relativity.
@@
Figure 1: Scheme of ET for the triangular configuration. Each detector ET1,2,3 is composed of an interferometer optimized for low-frequency detection and another optimized for high-frequency detection for a total of six.
Since 2014, the Sos Enattos [59, 60, 61, 63, 62, 64] and EMR [65, 66, 67, 68] sites have been the subject of characterization studies aimed at assessing their suitability to host ET. However, except Ref. [63] and the recent work published by Ref. [69], no works are assessing the impact of site dependent noise on the detectability of GW sources. Therefore, the scope of this paper is to provide a first estimation of how site dependent low-frequency noise may affect the SNR of CBC signals after modifying the ET sensitivity curve according to the actual seismic noise measured at each of the aforementioned ET candidate sites. We think that this approach may provide some relevant information for both the site selection process and the design of noise suppression systems for ET. In particular, extending the methodology already shown in Refs.[63] and [70], we investigate the case of IMBH and BNS focusing on the $[2,10]\rm\,Hz$ range, where ambient noise is more prominent [61, 62, 63] and may affect early warnings. Beyond 10 Hz, noise contributions from the detector hardware dominate the noise budget and, therefore, are out of the scope of our work. As a consequence of the focus in the low frequency range, our goal is not to claim whether a merger will be detected or not (merger frequencies are way beyond the considered frequency range), but only if it will be possible to observe the inspiral preceding the merger for multimessenger purposes. On the other hand, the details of how parameter estimation and other cosmological implications would be affected are beyond the goal of this work and will not be discussed here, but may represent one of the future developments of this study.
The paper is organized in the following way: in Section 2 we present the method; Section 3 talks through the data used for this study; the results are discussed in Sections 4.1 and 4.2. Finally, the conclusions are debated in Sections 5 and 6.
@@
2 Method
@@
To estimate the impact of site-dependent noise over the SNR of CBC events, we first calculate how site noise changes the ET design sensitivity curve. For this study we use the design curve already applied for Ref. [8]. Depending on the frequency band, different contributions affect the ET noise budget curve like seismic, Newtonian, thermal, and quantum noise. Among these noise sources, the dominant contribution in the frequency region between 2 and 10 Hz is the NN. Beyond 10 Hz, other contributions, like quantum and thermal noise from the detector hardware, dominate the noise budget [80, 81, 57, 82]. NN is caused by gravity fluctuations from mass distributions that change around the detector with time. These fluctuations are generated by dynamic variations of atmospheric and ground densities in the vicinity of the test masses of the detectors. Ground density variations are caused by ground motion, i.e., seismic waves, which are the only contribution to NN considered in this work. Atmospheric NN is not taken into account due to the current lack of dedicated environmental sensors installed at the ET candidate sites.
@@
which is Equation 5 of [63] and Equation 6 of [82], where $\vec{\xi}_{P,S}$ is the seismic displacement vector for the compressional (P) and shear (S) waves, respectively, and $\omega$ is the angular frequency. Equation 1 holds for a spherical cavern in an infinite homogeneous medium as it is clearly stated after Equation 6 of Ref.[82]. An even more detailed derivation can be found in Section 3.3 of Ref. [57]. In general, equation 1 is evaluated assuming:
@@
The ASD of seismic noise ($\tilde{x}$ of Equation 1) is evaluated calculating single fast Fourier tranforms (FFT) of the horizontal components of seismic noise, which, contrary to the vertical component, are dominated by body waves waves [83, 84], on $120\rm\,s$ windows. The length of the time window is chosen according to the typical maximum duration of an IMBH signal (Figure 4(b)) and on the time segment duration proposed for CBC multiband analysis for BNS [20]. Each FFT segment is then used to evaluate the probabilistic power spectral density (PPSD) of seismic noise over 2 years of data from which we infer the percentiles (10th, 50th, and 90th) of the spectral distributions. Under these assumptions, the ET noise budget that we use takes into account the NN evaluated from the seismic noise of each candidate site for the 10th, 50th, and 90th percentile of seismic spectra. Since a simple NN projection alone does not show the effect of non-stationarity of NN (from now on defined glitchness) we also use a more effective indicator, called the Noise to Target Ratio (NTR) and derived from Ref. [63] as:
$\rm{NTR}=\sqrt{\frac{1}{\Delta f}\int_{f_{1}}^{f_{2}}{df\frac{S_{n}(f)}{S_{n,\rm{ET}}(f)}}},$ (3)
@@
After inferring the modified ET sensitivity curves, we gather the events of interest for our study from astrophysical catalogs and we use PyCBC [85] to generate and inject signals in simulated ET noise. Waveforms are generated using the IMRPhenomD [76, 77] GW approximant for IMBH and IMRPhenomPv2_NRTidalv2 [78] for BNS, both available in the LIGO-Virgo-KAGRA Algorithm Library (LAL) [86]. To take into account the antenna pattern of the detector, the ET triangle is simulated considering three properly oriented L-shaped co-located detectors, with an arm length of $10\rm\,km$ and an arm angle of $60^{\circ}$. The variation of the antenna pattern with time is also taken into account. Since the actual orientation of the ET triangle has not been decided yet, we choose an arbitrary positioning of the detector. For the Sardinia candidate site, ET is assumed to be located at the Sos Enattos mine at $[40.44^{\circ}\rm\,N,9.44^{\circ}\rm\,E]$, whereas for the EMR site, we have chosen the village of Terziet at $[50.73^{\circ}\rm\,N,5.91^{\circ}\rm\,E]$, a proposed location for at least one of the ET vertexes, as a reference. For each detector, noise is obtained by generating a noise time series according to the given sensitivity curve and the same sources are taken into account. Waveforms are injected into the noise of each detector composing ET (labeled ET1, ET2 and ET3 respectively).
@@
where $\tilde{s}(f)$ is the Fourier Transform (FT) of the data, $\tilde{h}^{*}(f)$ is the FT of the template and $h_{cos}$/$h_{sin}$ are the two orthogonal phase components of the template shown in Equations 5 and 6 [87, 88]. In our case, $f_{\rm{min}}=2\rm\,Hz$ and $f_{\rm{max}}=10\rm\,Hz$ because, as mentioned above, we consider ambient noise only which modifies the ET sensitivity curve not beyond the $[2,10]\rm\,Hz$ range. Moreover, taking into account Equations 7 and 8 we note that MSNR out of this bandwidth gives the same contribution to the total MSNR in all cases and would not bring relevant information. Therefore we restrict the calculation between 2 Hz and 10 Hz only. We also point out that we assume perfect knowledge of the source, i.e., matched filter is calculated using the same waveform template injected in the data.
3 Data
3.1 Seismic